The joint pmf of X and Y is defined by f(x,y)=x + 2y 18, x=1,2; y=1,2

(a) Find Cov(X,Y).

(b)Find E(X|Y=1)
Answer #1

a)

E(*X) =\sum\sum x*P(x,y) =\sum\sum(x2+2xy)/18

=(1^2+2*1*1)/18+(1^2+2*1*2)/18+(2^2+2*2*1)/18+(2^2+2*2*2)/18 =14/9

E(Y) =\sum\sum y*P(x,y) =\sum\sum(yx+2y2)/18

=(1*1+2*1^2)/18+(1*2+2*2^2)/18+(2*1+2*1^1)/18+(2*2+2*2^2)/18 =29/18

E(XY) =\sum\sum xy*P(x,y) =\sum\sum(x2y+2xy2)/18 =2.5

Covar(y,x)=E(XY)-E(X)*E(Y)= 2.5-(14/9)*(29/18) =-1/162

b)P(Y=1) =P(X=1,Y=1)+P(X=2,Y=1) =(1+2*1)/18+(2+2*1)/18 =7/18

E(X|Y=1) =\sumxP(x|y=1)

=(18/7)*(1*3/18+2*4/18)=11/7